OpenStudy (anonymous):

Can someone show me how to do this problem step by step? (8+9i)(8-9i)

OpenStudy (lalaly):

ok expand the brackets the same way u do when u have x\[(8+9i)(8-9i)\]\[64-72i+72i -9i^2\]right?

OpenStudy (lalaly):

so the -72i +72i =0 so u cancel them and do u know what i^2 =?

OpenStudy (lalaly):

\[i=\sqrt{-1} \]\[i^2= -1\]

OpenStudy (lalaly):

so it becomes \[64-9(-1) = 64+9\]

OpenStudy (lalaly):

=73

OpenStudy (anonymous):

64+81=145

OpenStudy (lalaly):

oops my bad\[64-72i+72i -81i^2 = 64-81(-1) = 64+81= 145\]

OpenStudy (anonymous):

thank you

OpenStudy (lalaly):

=)